OpenStudy (anonymous):

price of stock fell from $36.88 to $27.65 find the % of decrease in price

OpenStudy (anonymous):

the % of decrease in price is equal to %= (36.88-27.65) / (original price = 36.88)

OpenStudy (anonymous):

36.88-27.65=9.23/36.88=3.99

OpenStudy (anonymous):

wrong 25.027115%

OpenStudy (anonymous):

you see what you have to do is divide 100 into 36.88 and then you find 36.88-27.65 giving you 9.23 then you multiply 9.23 by the result of your division of 100 into 36.88

OpenStudy (anonymous):

36.88-27.65=9.23/36.88=0.2502x100%=25.02%

OpenStudy (anonymous):

Exactly! Great job! :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Anytime :)